Calculate the energy carried by an x-rays of wavelength 7.5×10-10m(h=6.6×10-34js, vorc=3×108m/s)(Hmt:E=hc)


1
Expert's answer
2021-10-27T08:48:57-0400

We can find the energy carried by an x-rays of wavelength "\lambda" as follows:


"E=\dfrac{hc}{\lambda}=\dfrac{6.626\times10^{-34}\ J\times s\times3.0\times10^8\ \dfrac{m}{s}}{7.5\times10^{-10}\ m}=2.65\times10^{-16}\ J."